OpenStudy (anonymous):

A van traveling down a slope with a uniform acceleration of 2.15 meters/second2 attains a speed of 20.00 meters/second after 7.00 seconds. What is the initial velocity of the van?

OpenStudy (anonymous):

\[v _{f}=v _{i}+a \times (t _{f}-t _{i})\] \[20 = v _{i}+2.15 \times (7-0)\] \[20=v _{i}+15.05\] \[v _{i} = 4.95_{m/s}\]